# Cyclotomic Factors and LRS-Degeneracy

### v3

@@ Abstract
[TUK]organization=Fachbereich Mathematik, RPTU-Kaiserslautern, addressline=Gottlieb-Daimler-Strasse, city=Kaiserslautern, postcode=D-67663, state=Rhineland-Palatinate, country=Germany
@@
[UPA]organization=Fakultät für Informatik und Mathematik, Universität Passau, addressline=Innstrasse 33, city=Passau, postcode=D-94032, state=Bavaria, country=Germany
@@ 1 Introduction
Our emphasis is on practical utility, so we do not present complexity analyses of the algorithms, since they are of little utility in this context (and would detract from the focus of this paper). We do present timings; in particular, we show that our LRS-degeneracy algorithm is significantly faster than the two which were presented in [3].
@@ 2 Notation, Terminology, Preliminaries
We shall often restrict attention to square-free, content-free polynomials in ${\mathbb{Z}}[x]$: a general polynomial in ${\mathbb{Q}}[x]$ may readily be rescaled by clearing denominators and dividing out the content to yield a content-free polynomial in ${\mathbb{Z}}[x]$. We can easily compute the radical (i.e. product of the factors in a square-free factorization) as $f/\gcd(f,f^{\prime})$ where $f^{\prime}$ denotes the formal derivative of $f$. We shall write the reverse of a non-zero polynomial as $\mathop{\rm rev}(f)\,=\,x^{d}\,f(\frac{1}{x})$ where $d=\deg(f)$.
@@
The Graeffe transform exists in the same ring as the original polynomial. Also for $k=2$ and $k=3$ the Graeffe transform can be computed quickly and easily: e.g. $G_{2}(f)\,=\,(-1)^{d}(f_{even}^{2}-x\,f_{odd}^{2})$ where $f_{even}$ and $f_{odd}$ are defined by $f(x)\,=\,f_{even}(x^{2})+x\,f_{odd}(x^{2})$.
@@
Let $k\in{\mathbb{N}}_{\geq 2}$ and $f\in{\mathbb{Z}}[x]$. Then we say that $f$ is $\bm{k}$-LRS-degenerate iff $f$ has two distinct roots $\alpha,\beta\in{\mathbb{C}}$ such that $\alpha/\beta$ is a primitive $k$-th root of unity. We also say simply that $f$ is LRS-degenerate to mean that there exists at least one order $k\in{\mathbb{N}}_{\geq 2}$ such that $f$ is $k$-LRS-degenerate.
@@
Let $f,g\in{\mathbb{C}}[x]$. Then we say that $f$ and $g$ are LRS-degeneracy equivalent iff there are scale factors $\lambda,\mu\in{\mathbb{C}}_{\neq 0}$ such that $f(x)=\mu\,g(\lambda x)\;$ or $f(x)=\mu\,x^{d}\,g(\lambda/x)$ for some $d\in{\mathbb{N}}$.
@@ 3.1.1 Coefficient of $x^{d-1}$ and the special case $f(x)=g(x^{r})$
    Otherwise we apply all the “square-free index” checks below to $g(x)=f(x^{1/r})$. If $g(x)$ is not cyclotomic then neither is $f(x)$. If however $g(x)$ is cyclotomic with index $k$, we then check whether $\mathop{\rm rad}(r)\mid k$. If so then $f(x)$ is cyclotomic with index $rk$, otherwise $f$ is not cyclotomic.
@@ 3.2 Testing for a cyclotomic polynomial with square-free index
Here we assume that all the checks above have been made (esp. steps Q4a-e), so that if $f$ is cyclotomic then its index must be square-free (and at least 3). As already indicated, the check we give here involves “smart brute force”, and is decidedly more computationally costly than the quick checks above.
@@
In Algorithm 1 (CycloIndex) in Step 17, it suffices to compute the $\frac{d}{2}$-prefix and then check that $f$ is palindromic.
@@
7:  Initialize $L_{new}\leftarrow[\,]$, an empty list
@@ 3.2.1 Computing an $m$-prefix
Let $n\in{\mathbb{Z}}_{>2}$ be square-free. Since cyclotomic polynomials are palindromic, prefixes and suffixes are equivalent. It is convenient here to discuss suffixes: namely $\Phi_{n}(x)\mod x^{m}$. We can compute an $m$-suffix cheaply via the Möbius inverted product:
$\Phi_{n}(x)\;=\;\prod_{d\mid n}(x^{d}-1)^{\mu(n/d)}$
Observe that multiplying by $x^{d}-1$ is just a shift-and-subtract operation, and indeed dividing by $x^{d}-1$ can also be implemented as a shift-and-add operation (working up from lowest to highest degree). We can also ignore all $d\geq m$.
So long as $m$ is not too large, all arithmetic can be done with machine integers: the outermost coefficients of a cyclotomic polynomial cannot be too large. A table of values is available as sequence A138474 at OEIS [8]. On a 32-bit computer overflow cannot occur with $m\leq 808$; and on a 64-bit computer all $m\leq 1000$ are safe — the table at OEIS stops at index 1000; the true limit for 64-bit computers is likely higher.
Also if $n$ has fewer than 4 odd prime factors then overflow cannot occur: if $n$ has 1 or 2 odd prime factors then the height is 1, and if there are 3 odd prime factors then the height is at most $2p/3$ where $p$ is the smallest odd prime factor [9].
In Algorithm 1 at Step 10 we use $f_{m}$, the $m$-prefix of $f$. Based on the sequence A138474, we incorporate a coefficient height check while generating $f_{m}$ from $f$, provided $m\leq 1000$: thus $f_{m}$ can be represented using machine integers (otherwise $f$ is surely not cyclotomic). Our implementation gives an error if ever $m>1000$ occurs.
@@ 3.3 Testing via evaluation
We describe an alternative approach which is a special case of our method for identifying cyclotomic factors presented in Section 4. Our implementation was not as fast as the “matching prefix” method described in Section 3.2, at least partly because in CoCoALib “reading” the coefficients of a polynomial is costly.
We reiterate the assumption that if $f$ is cyclotomic then its index is square-free. We can perform the following further “quick checks” (which can be performed together in a single scan over the coefficients):
  • Q5a
  • Q5b
    Check that $f$ is palindromic.
  • Q5c
  • Q5d
3.3.1 Obtaining a candidate index (via divisibility)
@@
We could make the check faster by computing the canonical remainder of $x^{r}\mod f(x)$ for $r=2d,\,3d,\,4d,\,\ldots$. Bounds on the inverse totient function show that for degrees <36 495 360<36{\,}495{\,}360 (see sequence A355667 at OEIS [8]) we can stop at $r=6d$ since $k$ cannot exceed this value. If the polynomial is indeed $\Phi_{k}$ then once we try the first exponent $r\geq k$ we obtain remainder $x^{s}$ (with s=rks=r-k), and it is easy to identify if the remainder has this form. From this remainder we easily obtain $k$ as rsr-s. However again, actually computing these polynomial remainders would still be rather costly.
Instead we mimic the above but under the mapping $x\mapsto 2$. We evaluate $f(2)$, and check that $2^{d-1}<f(2)<2^{d+1}$. Then we compute $2^{r}\mod f(2)$ for $r=2d,\,3d,\,4d,\,\ldots,7d$; to advance to the next value in the sequence we just multiply by $2^{d}$ then reduce modulo $f(2)$. Observe that on a binary computer it is easy to detect if the remainder is of the form $2^{s}$, in which case we obtain the candidate index k=rsk=r-s. We finally verify that $\deg(f)=\phi(k)$ and if so, also that $f=\Phi_{k}$ by computing $\Phi_{k}$.
@@
We note that the final check (that $f=\Phi_{k}$) is necessary, at least for higher degrees, and index $k$ not of the form $p$ or $2p$ for some prime $p$. Consider the polynomial $\Phi_{k}(x)+x^{d/2-3}\,(x^{2}-1)^{2}\,(2x^{2}-5x+2)$ where $d=\deg(\Phi_{k})$; this is palindromic, and will pass all the evaluation tests, since the added polynomial vanishes at $x=\pm 1$ and $x=2$. For higher degrees the coefficient height check will pass also.
@@ 3.3.2 Supporting arguments for evaluation at 2
We justify the algorithm outlined in Section 3.3.
@@ 3.3.4 Obtaining a candidate index (via table-lookup)
Let $S$ be the set of cyclotomic polynomials of degree up to $1000$ and having square-free index. Let the prime $p=57737$. Then the mapping $S\to{\mathbb{F}}_{p}$ sending $f\mapsto f(2)\mod p$ is 1–1. So we can precompute a table of indexes: e.g. an array whose $r$-th entry is the index $k$ such that $r=\Phi_{k}(2)\mod p$. We fill “empty” array entries with a value which is not a valid index (e.g. 0).
We can use the table as follows: if $\deg(f)\leq 1000$, compute $r=f(2)\mod p$. If the table entry corresponding to residue $r$ is “empty” then $f$ is surely not cyclotomic. Otherwise, let $k$ be the table entry, and verify that $\phi(k)=\deg(f)$ and that $\mu(k)$ is equal to the coefficient of $x^{d-1}$, if so then maybe check that $2^{k}\equiv 1\pmod{f(2)}$, and finally compute $\Phi_{k}$ then test whether $f=\Phi_{k}$.
@@ 3.4 Timings
Here we give some sample timings of our implementation of the “prefix method” (from Section 3.2) in CoCoALib. We restrict to the case of identifying the index of a cyclotomic polynomial without verification in Step 17, because the verification step involves computing the full polynomial, and this verification dominates. To present meaningful timings we consider only “difficult” cases where a large prefix length is needed, and there are many preimages to consider: for these examples, the smaller factor of each index is a lower bound on the prefix length required.

### v4

@@ Abstract
[TUK]organization=Fachbereich Mathematik, RPTU University Kaiserslautern–Landau, organizationsep= , addressline=Gottlieb-Daimler-Strasse, city=Kaiserslautern, postcode=D-67663, country=Germany
@@
[UPA]organization=Fakultät für Informatik und Mathematik, Universität Passau, addressline=Innstrasse, city=Passau, postcode=D-94032, country=Germany
@@ 1 Introduction
Our emphasis is on practical utility, so we do not present asymptotic complexity analyses of the algorithms, since they are of little utility in this context (and would detract from the focus of this paper). We do present timings; in particular, we show that our LRS-degeneracy algorithm is significantly faster than the two which were presented in [9].
@@ 2 Notation, Terminology, Preliminaries
We shall often restrict attention to square-free, content-free polynomials in ${\mathbb{Z}}[x]$: a general non-zero polynomial in ${\mathbb{Q}}[x]$ may readily be rescaled by clearing denominators and dividing out the content to yield a content-free polynomial in ${\mathbb{Z}}[x]$. We can easily compute the radical (i.e. product of the factors in a square-free factorization) as $f/\gcd(f,f^{\prime})$ where $f^{\prime}$ denotes the formal derivative of $f$. We shall write the reverse of a non-zero polynomial as $\mathop{\rm rev}(f)\,=\,x^{d}\,f(\frac{1}{x})$ where $d=\deg(f)$.
@@
Definition 2.3.
Let $f=\sum_{j=0}^{d}a_{j}x^{j}\in{\mathbb{C}}[x]$. We define the height of $f$ to be $H(f)\;=\;\max\{|a_{0}|,|a_{1}|,\ldots,|a_{d}|\}$ with the special case $H(0)=0$.
@@
The Graeffe transform exists in the same ring as the original polynomial. Also for $k=2$ and $k=3$ the Graeffe transform can be computed quickly and easily: e.g. $G_{2}(f)\,=\,(-1)^{d}(f_{even}^{2}-x\,f_{odd}^{2})$ where $d=\deg(f)$, and $f_{even}$ and $f_{odd}$ are defined by $f(x)\,=\,f_{even}(x^{2})+x\,f_{odd}(x^{2})$.
@@
Let $k\in{\mathbb{N}}_{\geq 2}$ and non-zero $f\in{\mathbb{C}}[x]$. Then we say that $f$ is $\bm{k}$-LRS-degenerate iff $f$ has two distinct roots $\alpha,\beta\in{\mathbb{C}}$ such that $\alpha/\beta$ is a primitive $k$-th root of unity. We also say simply that $f$ is LRS-degenerate to mean that there exists at least one order $k\in{\mathbb{N}}_{\geq 2}$ such that $f$ is $k$-LRS-degenerate. We are primarily interested in the case $f\in{\mathbb{Z}}[x]$.
@@
Let non-zero $f,g\in{\mathbb{C}}[x]$. Then we say that $f$ and $g$ are LRS-degeneracy equivalent iff there are scale factors $\lambda,\mu\in{\mathbb{C}}_{\neq 0}$ such that $f(x)=\mu\,x^{d}\,g(\lambda x)\;$ or $f(x)=\mu\,x^{d}\,g(\lambda/x)$ for some $d\in{\mathbb{Z}}$. It is elementary to show that this is an equivalence relation.
@@ 3.1.1 Coefficient of $x^{d-1}$ and the special case $f(x)=g(x^{r})$
    Otherwise we conduct the “square-free index” test of Section 3.3 on $g(x)=f(x^{1/r})$. If $g(x)$ is not cyclotomic then neither is $f(x)$. If however $g(x)$ is cyclotomic with index $k$, we then check whether $\mathop{\rm rad}(r)\mid k$. If so then $f(x)$ is cyclotomic with index $rk$, otherwise $f$ is not cyclotomic.
@@
3.2 Further quick checks
We present some further simple checks; unlike the checks Q1–Q4b, these checks can access all the coefficients of the polynomial, so they are typically costlier. Checks Q5c and Q5d are valid only for cyclotomic polynomials with square-free index, so they must be conducted only after all checks in Section 3.1 have been made; in contrast, checks Q5a and Q5b are independent of the other checks.
  • Q5a
    Check that $f$ is palindromic.
  • Q5b
  • Q5c
  • Q5d
Observe that all four of these checks can be performed together in a single scan over the coefficients. Also, checks Q5c and Q5d can be seen as a special case of our method for identifying cyclotomic factors presented in Section 4.
Remark 3.2.
We implemented these checks in CoCoALib, but found that they are not as fast as the “matching prefix” method described in Section 3.3. This is partly explained by the fact that in CoCoALib “reading” all the coefficients of a polynomial is costly. Thus our final implementation does not actually employ these checks.
@@ 3.3 Testing for a cyclotomic polynomial with square-free index
Here we assume that all the checks from Section 3.1 have been made, so that if $f$ is cyclotomic then its index must be square-free (and at least 3). As already indicated, the test we give here involves “smart brute force”, and is decidedly more computationally costly than the quick checks in Section 3.1.
@@
In Algorithm 1 (CycloIndex) in Steps 16 and 17, it suffices to verify the $\frac{d}{2}$-prefix and then check that $f$ is palindromic.
@@
7:  $L_{new}\leftarrow[\,]$, an empty list
@@
Remark 3.5.
A referee described a more efficient version of Algorithm 1 (CycloIndex): an outer loop iterates over the candidate indexes, and an inner loop checks each candidate by constructing successively longer prefixes until the candidate is either excluded or confirmed as sole candidate for final verification. This approach will be faster if the input polynomial is indeed cyclotomic, and its index appears early in the list of candidates. The difference in overall speed is likely minor since the final verification is typically the main part of the total computational cost. Moreover, Algorithm 1 (CycloIndex) as presented can be trivially modified to skip the final verification (Steps 16 and 17): it then determines the index when the input is known to be cyclotomic — see Section 3.5. It is not clear whether the referee’s variant could easily be adapted to forego the final verification.
@@ 3.3.1 Computing an $m$-prefix
Let $n\in{\mathbb{Z}}_{>2}$ be square-free. Since cyclotomic polynomials are palindromic, prefixes and suffixes are equivalent. It is convenient here to discuss suffixes: namely $\Phi_{n}(x)\bmod x^{m}$. We can compute an $m$-suffix cheaply via the Möbius inverted product:
$\Phi_{n}(x)\;=\;\prod_{d\mid n}(x^{d}-1)^{\mu(n/d)}\;=\;\prod_{d\mid n}(1-x^{d})^{\mu(n/d)}$
We may negate the factors because $n$ has an even number of divisors. Observe that multiplying by $1-x^{d}$ is just a shift-and-subtract operation, and indeed dividing by $1-x^{d}$ can also be implemented as a shift-and-add operation (working up from lowest to highest degree). We may ignore all $d\geq m$.
3.3.2 Coefficient Bounds and Machine Integers
In Algorithm 1 (CycloIndex) at Step 10 we use $f_{m}$, the $m$-prefix of $f$. We incorporate a coefficient height check while generating $f_{m}$ from $f$; if a coefficient is too large then we know that $f$ cannot be cyclotomic. Furthermore, so long as $m$ is not too large, all arithmetic can be done with machine integers: the outermost coefficients of a cyclotomic polynomial cannot be too large. A table of bounds for each of the outermost coefficients is available as sequence A138474 at OEIS [16]. On a 32-bit computer overflow cannot occur with $m\leq 808$; and on a 64-bit computer all $m\leq 1\,000$ are safe — the table at OEIS stops at index 1 0001\,000; the true limit for 64-bit computers is likely considerably higher.
For prefixes of length m>1 000m>1\,000 we can use the tabulated values of height bounds for cyclotomic polynomials (up to index 2 583 303 5552\,583\,303\,555 — see Section 3.4.4). Additionally, for each candidate index we can apply the formulas below to obtain potentially better bounds (especially when the index has few prime factors). Our implementation gives an error if the best bound found exceeds the largest machine integer.
In [8] there are formulas bounding the height of $\Phi_{n}$ and of its power-series inverse $1/\Phi_{n}(x)$ for an odd, square-free index $n=p_{1}p_{2}\cdots p_{k}$ where $p_{1}<p_{2}<\cdots<p_{k}$. Interestingly, all the bounds are independent of the largest two prime factors. If $k<3$ then the height bound is $1$. If $k=3$ then the bound is ${\textstyle\frac{2}{3}}p_{1}$ according to preprint [12], and on the basis of this result the article [8] establishes bounds on the height of $\Phi_{n}$ for $k=4,5,6$:
$\frac{2}{3}p_{1}^{3}p_{2}\qquad\frac{2}{9}p_{1}^{7}p_{2}^{3}p_{3}\qquad\frac{32}{729}p_{1}^{15}p_{2}^{7}p_{3}^{3}p_{4}$
The bounds on the height of the inverse have a similar form but with different constant factors. If the proof in [12] is not confirmed as correct then [8] obtains other constant factors. A fully general formula is given in [6]. So far we have implemented just the simple cases $k=1$ and $k=2$.
If some of the prime factors of $n$ are greater than the prefix length $m$, we may sometimes obtain lower bounds. For instance, if $n=p_{1}p_{2}p_{3}$ and $p_{1}<m\leq p_{2}$ then the only non-trivial truncated factors in the Möbius product are the same as those required to compute the $m$-prefix of $\Phi_{p_{1}}$. Thus the height bound for index $p_{1}$ limits the $m$ outermost coefficients also of $\Phi_{n}$; indeed, in this instance we know that the coefficients of $x^{e}$ for $p_{1}\leq e<p_{2}$ must be zero. Depending on whether the number of prime factors bigger than $m$ is even or odd, we use the height bound for the cyclotomic polynomial or for its inverse.
@@ 3.4 Testing via evaluation
We recall the quick checks Q5c and Q5d from Section 3.2: these perform evaluation at $1$ and at 1-1 respectively, combined with a check that the resulting value is permissible. We now show that evaluation at $2$ provides us with a simple way of obtaining a unique candidate index (or proof that the polynomial is not cyclotomic). However, our implementation of the prefix method of Algorithm 1 (CycloIndex) turned out to be faster.
3.4.1 Obtaining a candidate index (via remainders of powers)
@@
We could make the check faster by computing the canonical remainder of $x^{r}\bmod f(x)$ for $r=2d,\,3d,\,4d,\,\ldots$. Bounds on the inverse totient function show that for degrees <36 495 360<36{\,}495{\,}360 (see sequence A355667 at OEIS [16]) we can stop at $r=6d$ since $k$ cannot exceed this value. If the polynomial is indeed $\Phi_{k}$ then once we try the first exponent $r\geq k$ we obtain remainder $x^{s}$ (with s=rks=r-k), and it is easy to identify if the remainder has this form. From this remainder we easily obtain $k$ as rsr-s. However again, actually computing these polynomial remainders would still be rather costly.
Instead we mimic the above but under the mapping $x\mapsto 2$. We evaluate $f(2)$, and check that $2^{d-1}<f(2)<2^{d+1}$. Then we compute $2^{r}\bmod f(2)$ for $r=2d,\,3d,\,4d,\,\ldots,7d$; to advance to the next value in the sequence we just multiply by $2^{d}$ then reduce modulo $f(2)$. Observe that on a binary computer it is easy to detect if the remainder is of the form $2^{s}$, in which case we obtain the candidate index k=rsk=r-s. We finally verify that $\deg(f)=\phi(k)$ and if so, also that $f=\Phi_{k}$ by computing $\Phi_{k}$.
@@
We note that the final check (that $f=\Phi_{k}$) is necessary, at least for higher degrees and index $k$ not of the form $p$ or $2p$ for some prime $p$. Consider the polynomial $\Phi_{k}(x)+x^{d/2-3}\,(x^{2}-1)^{2}\,(2x^{2}-5x+2)$ where $d=\deg(\Phi_{k})$; it is palindromic, and will pass all the evaluation tests, since the added polynomial vanishes at $x=\pm 1$ and $x=2$. For higher degrees the coefficient height check will pass also.
3.4.2 Explicit Algorithm
The algorithm has to exclude just a few edge cases: namely, $\Phi_{1}$ and $\Phi_{6}$ which are exceptions in Zsigmondy’s theorem. It is convenient, but not necessary, to exclude $\Phi_{2}$ as well. We assume that the $\bmod$ function returns the least non-negative remainder.
We note that on a binary computer it is quick and simple to recognize if an integer is a power of $2$ and to determine the corresponding exponent: we do this in Algorithm 2 (CycloIndexByEval) Step 8.
Algorithm 2 (CycloIndexByEval)
[…]
2:
[…]
4:if $v\leq 2^{d-1}$ or $v\geq 2^{d+1}$ then return not cyclotomic end if
[…]
6:for $r^{\prime}\leftarrow 2$ to 77 do $\triangleright$ where $r^{\prime}\leftrightarrow{\textstyle\frac{r}{d}}$
7:  $a\leftarrow(ap)\bmod v$
8:  if $a=2^{s}$ for some $s$ then return $r^{\prime}d-s$ end if
9:end for
10:return not cyclotomic
@@ 3.4.3 Supporting arguments for evaluation at 2
We justify Algorithm 2 (CycloIndexByEval) presented in Section 3.4.2.
@@ 3.4.5 Obtaining a candidate index (via table-lookup)
Let $S$ be the set of cyclotomic polynomials of degree up to $1000$ and having square-free index. Let the prime $p=57737$. Then the mapping $S\to{\mathbb{F}}_{p}$ sending $f\mapsto f(2)\bmod p$ is 1–1. So we can precompute a table of indexes: e.g. an array whose $r$-th entry is the index $k$ such that $r=\Phi_{k}(2)\mod p$. We fill “empty” array entries with a value which is not a valid index (e.g. 0).
We can use the table as follows: if $\deg(f)\leq 1000$, compute $r=f(2)\bmod p$. If the table entry corresponding to residue $r$ is “empty” then $f$ is surely not cyclotomic. Otherwise, let $k$ be the table entry, and verify that $\phi(k)=\deg(f)$ and that $\mu(k)$ is equal to the coefficient of $x^{d-1}$; if so then maybe check that $2^{k}\equiv 1\pmod{f(2)}$, and finally compute $\Phi_{k}$ then test whether $f=\Phi_{k}$.
@@ 3.5 Timings
Here we give some sample timings of our implementation of the “prefix method” (from Section 3.3) in CoCoALib. We restrict to the case of identifying the index of a cyclotomic polynomial without verification in Steps 16 and 17, because the verification step involves computing the full polynomial, and this verification dominates. To present meaningful timings we consider only “difficult” cases where a large prefix length is needed, and there are many preimages to consider: for these examples, the smaller factor of each index is a lower bound on the prefix length required.
@@ 4.1.1 Zsigmondy’s Theorem
Zsigmondy’s theorem [18] tells us that if we have the sequence $s_{n}=a^{n}-b^{n}$ for coprime integers $a>b>0$ then for every index the element $s_{k}$ has a primitive prime. The only exceptions are: